OpenStudy (anonymous):

What are the domain and range of f(x)=log(x-1)+2? I WILL GIVE MEDALS AND WHATEVER YOU WANT

OpenStudy (anonymous):

\(\$50\) ?

OpenStudy (anonymous):

range first , range of the log is \((-\infty, \infty)\) or \(\mathbb{R}\)

OpenStudy (anonymous):

Lol @satellite73

OpenStudy (anonymous):

domain, solve \[x-1>0\] get \[x>1\] in one step domain \(x>1\) or \((1,\infty)\)

OpenStudy (anonymous):

for another $50 i will guarantee those answers are correct

OpenStudy (anonymous):

guess they will have to pay me when they get back on line

OpenStudy (anonymous):

Thank you so much if I had $50 I'd give it to you